Question:
The perimeter of a rectangular field is 60 meters. Its area is 200 meters. Find its dimensions.

Please explain what you do.
Thanks

Answer by checkley79(3341)

AREA=LW
200=LW
L=200/W
P=2L+2W
60=2(200/W)+2W
60=400/W+2W/1 COMBINE THE 2 TERMS ON THE RIGHT BY FINDING
A COMMON DENOMINATOR. IN THIS CASE IT IS W.
60=(400+2W*W)/W
60=(400+2W^2)/W CROSS MULTIPLY.
60W=400+2W^2
2W^2-60W+400=0
2(W^2-30W+200)=0
2(W-20)(W-10)=0
W-20=0
W=20 ANS. L=200/20=10 ANS.
W-10=0
W=10 ANS. L=200/10=20 ANS.
200=10*20
200=200